OpenStudy (anonymous):

Select the expression equal to √48x^4y^4

OpenStudy (anonymous):

\(\large \sqrt{48x^4y^4}\) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what expression equals to that

OpenStudy (anonymous):

\(\large \sqrt{48x^4y^4}=\sqrt{(3\times4^2)(x^2)^2(y^2)^2}=4x^2y^2\sqrt{3}\)

OpenStudy (anonymous):

thank you!!! :)

OpenStudy (anonymous):

^_^ \(\Large {\color{blue}{Ur\ \ Welcome}}\)